### v1

# Vision-and-Language Navigation Today and Tomorrow: A Survey in the Era of Foundation Models

Abstract
Vision-and-Language Navigation (VLN) has gained increasing attention over recent years and many approaches have emerged to advance their development. The remarkable achievements of foundation models have shaped the challenges and proposed methods for VLN research. In this survey, we provide a top-down review that adopts a principled framework for embodied planning and reasoning, and emphasizes the current methods and future opportunities leveraging foundation models to address VLN challenges. We hope our in-depth discussions could provide valuable resources and insights: on one hand, to milestone the progress and explore opportunities and potential roles for foundation models in this field, and on the other, to organize different challenges and solutions in VLN to foundation model researchers.
Vision-and-Language Navigation Today and Tomorrow:
A Survey in the Era of Foundation Models
Yue Zhang1thanks: Authors contributed equally to this work.  Ziqiao Ma211footnotemark: 1  Jialu Li311footnotemark: 1  Yanyuan Qiao411footnotemark: 1  Zun Wang311footnotemark: 1 Joyce Chai2thanks: Supervisors contributed equal advising to this work.  Qi Wu422footnotemark: 2  Mohit Bansal322footnotemark: 2  Parisa Kordjamshidi122footnotemark: 2 1Michigan State University  2University of Michigan 3University of North Carolina at Chapel Hill  4University of Adelaide
1 Introduction
Developing embodied agents that are capable of interacting with humans and their surrounding environments is one of the longstanding goals of Artificial Intelligence (AI) (Nguyen et al., 2021; Duan et al., 2022). These AI systems hold immense potential for real-world applications to serve as multi-functional assistants in daily life, such as household robots (Szot et al., 2021), self-driving cars (Hu et al., 2023), and personal assistants (Chu et al., 2023). One formal problem setting to advance this research direction is Vision-and-Language Navigation (VLN) (Anderson et al., 2018), a multimodal and cooperative task that requires the agent to follow human instructions, explore 3D environments, and engage in situated communications under various forms of ambiguity. Over the years, VLN has been explored in both photorealistic simulators (Chang et al., 2017; Savva et al., 2019; Xia et al., 2018) and real environments (Mirowski et al., 2018; Banerjee et al., 2021), leading to a number of benchmarks (Anderson et al., 2018; Ku et al., 2020; Krantz et al., 2020) that each presents slightly different problem formulations.
Figure 1: We organize challenges and solutions in VLN using the LAW framework (Hu and Shu, 2023). Specifically, the world model is a mental representation that encodes the visual perception and predicts the outcomes of actions. The human model is a mental representation that interprets the human intentions from textual instructions and contexts. They are a part of a VLN agent which reasons over multimodal input, engages in dialogue and plans the navigation actions accordingly.
Recently, foundation models (Bommasani et al., 2021), ranging from early pre-trained models like BERT (Kenton and Toutanova, 2019) to contemporary large language models (LLMs) and vision-language models (VLMs) (Achiam et al., 2023; Radford et al., 2021), have exhibited exceptional abilities in multimodal comprehension, reasoning, and cross-domain generalization. These models are pre-trained on massive data, such as text, images, audio, and video, and could further be adapted for a broad range of specific applications, including embodied AI tasks (Xu et al., 2024). Integrating these foundation models into VLN task marks a pivotal recent advancement for embodied AI research, demonstrated through significant performance improvements (Chen et al., 2021b; Wang et al., 2023f; Zhou et al., 2024a). Foundation models have also brought new opportunities to the VLN field, such as expanding the research focus from multi-modal attention learning and strategy policy learning to pre-training generic vision and language representations, hence enabling task planning, commonsense reasoning, as well as generalize to realistic environments.
Despite the recent impact of foundation models on VLN research, the previous surveys on VLN (Gu et al., 2022; Park and Kim, 2023; Wu et al., 2024) are from the pre-foundation-model era and mainly focus on the VLN benchmarks and conventional approaches, i.e., they are missing a comprehensive overview of the existing methods and opportunities leveraging foundation models to address VLN challenges. Especially with the emergence of LLMs, to the best of our knowledge, no review has yet discussed their applications in VLN tasks. Moreover, unlike previous efforts that discuss the VLN task as an isolated downstream task, the objective of this survey is twofold: first, to milestone the progress and explore opportunities and potential roles for foundation models in this field; second, to organize different challenges and solutions in VLN to foundation model researchers within a systematic framework. To build this connection, we adopt the LAW framework (Hu and Shu, 2023), where foundation models serve as backbones of world model and agent model. This framework offers a general landscape of reasoning and planning in foundation models, and is closely scoped with the core challenges in VLN.
Specifically, at each navigation step, the AI agents perceive the visual environment, receive language instructions from humans, and reason upon their representation of the world and humans to plan actions and efficiently complete navigation tasks. As shown in Figure 1, a world model is an abstraction that agents maintain to understand the external environment around them and how their actions change the world state (Ha and Schmidhuber, 2018; Koh et al., 2021). This model is part of a broader agent model, which also incorporates a human model that interprets the instructions of its human partner, thereby informing the agent’s goals (Andreas, 2022; Ma et al., 2023). To review the growing body of work in VLN and to understand the milestones achieved, we adopt a top-down approach to survey the field, focusing on fundamental challenges from three perspectives:
  • Learning a world model to represent the visual environment and generalize to unseen ones.
  • Learning a human model to effectively interpret human intentions from grounded instructions.
  • Learning a VLN agent that leverages its world and human model to ground language, communicate, reason, and plan, enabling it to navigate environments as instructed.
We present a hierarchical and fine-grained taxonomy in Figure 2 to discuss challenges, solutions, and future directions based on foundation models for each model. To organize this survey, we start with a brief overview of the background and related research efforts as well as the available benchmarks in this field (§2). We structure the review around how the proposed methods have addressed the three key challenges described above: world model (§3), human model (§4), and VLN agent (§5). Finally, we discuss the current challenges and future research opportunities, particularly in light of the rise of foundation models (§6).
2 Background and Task Formulations
In this section, we discuss the background, clarify the scope of this survey, define the VLN problem, and briefly overview the benchmarks.
Name World Human VLN Agent Dataset
Domain Environment Turn Format Gran. Type Act. Sp. Other Text Route
LANI/CHAI Misra et al. (2018) Indoors CHALET Single Multi Instr A - Disc Mani H H
R2R Anderson et al. (2018) Indoors Matterport3D Single Multi Instr A Robot Graph H P
R4R Jain et al. (2019) Indoors Matterport3D Single Multi Instr A Robot Graph H P
RxR Ku et al. (2020) Indoors Matterport3D Single Multi Instr A Robot Graph H P
SOON Zhu et al. (2021a) Indoors Matterport3D Single Multi Instr G Robot Graph H P
REVERIE Qi et al. (2020b) Indoors Matterport3D Single Multi Instr A, G Robot Graph Detect H P
VNLA Nguyen et al. (2019) Indoors Matterport3D Multi Multi Instr A, G Robot Graph T P
HANNA Nguyen and Daumé III (2019) Indoors Matterport3D Multi Multi Instr A, G Robot Graph H P
CVDN Thomason et al. (2020) Indoors Matterport3D Multi Restricted A Robot Graph H H
VLN-CE Krantz et al. (2020) Indoors Habitat, Matterport3D Single Multi Instr A Robot Disc H P
Robo-VLN Irshad et al. (2021) Indoors Habitat, Matterport3D Single Multi Instr A Robot Cont H P
RobotSlang Banerjee et al. (2021) Indoors Real Multi Freeform A Robot Disc H P
ALFRED Shridhar et al. (2020) Indoors AI2-THOR Single Multi Instr. A, G Robot Disc Mani H P
TEACh Padmakumar et al. (2022) Indoors AI2-THOR Multi Freeform A, G Robot Disc Mani H H
DialFRED Gao et al. (2022) Indoors AI2-THOR Multi Restricted A, G Robot Disc Mani H, T P
TouchDown Chen et al. (2019) Outdoors Google Street View Single Multi Instr A - Graph H P
Street Nav Hermann et al. (2020) Outdoors Google Street View Multi Multi Instr A - Disc T P
Talk2Nav Vasudevan et al. (2021) Outdoors Google Street View Single Multi Instr A, G - Disc H P
TtW De Vries et al. (2018) Outdoors Real Multi Freeform A, G - Disc H H
LCSD Sriram et al. (2019) Outdoors CARLA Single Multi Instr A Driving Disc H P
CDNLI Roh et al. (2020) Outdoors CARLA Multi Multi Instr A, G Driving Cont H, T H
SDN Ma et al. (2022) Outdoors CARLA Multi Freeform A, G Driving Disc, Cont H H
AerialVLN Liu et al. (2023b) Outdoors AirSim Single Multi Instr A, G Aerial Disc H H
ANDH Fan et al. (2023a) Outdoors xView Multi Freeform A, G Aerial Disc H H
Table 1: A summary of existing VLN benchmarks, taxonomized based on several key aspects: the world in which navigation occurs, the type of human interaction involved, the action space and tasks assigned to the VLN agent, and the methods of dataset collection. For the world, we consider their domain (either indoors or outdoors) and the environment. For the human, we consider their turns of interaction (either single or multiple turn), the format of communication (freeform dialogue, restricted dialogue, or multiple instructions), and the language granularity (action-directed and goal-directed). For the VLN agent, we consider their agent types (e.g., household robot, autonomous driving vehicles, or autonomous aerial vehicles), their action space (graph-based, discrete or continuous), and other additional tasks (manipulation and object detection). For dataset collection, we consider the text collection (by human or templated) and the route demonstrations (by human or planner).
2.1 Cognitive Underpinnings of VLN
Humans and other navigational animals demonstrate early understanding and strategies for navigating their environments (Rodrigo, 2002; Brand et al., 2015; Lingwood et al., 2018). For example, Gallistel (1990) describes two basic mechanisms: piloting, which involves environmental landmarks and computes distances and angles; and path integration, which calculates displacement and orientation changes through self-motion sensing. Central to understanding spatial navigation is the cognitive map hypothesis, suggesting that the brain forms a unified spatial representation to support memory and guide navigation (Epstein et al., 2017; Bellmund et al., 2018). For instance, Tolman (1948) observed that rats could adopt the correct novel path when familiar paths are blocked and landmarks are absent. Neuroscientists also discovered hippocampal place cells, indicating a spatial coordinate system that encodes landmarks and goals allocentrically (O’Keefe and Dostrovsky, 1971; O’keefe and Nadel, 1978). Recent studies also propose non-Euclidean representations, e.g., cognitive graphs, which illustrate the complexity of how we represent spatial knowledge of the world (Warren, 2019; Ericson and Warren, 2020). While visual and auditory perceptions are obviously integral to spatial representation (Klatzky et al., 2006), our linguistic skills and spatial cognition are also closely intertwined (Pruden et al., 2011). For instance, researchers have shown that understanding different aspects of spatial language can help with space-related tasks (Pyers et al., 2010), and that language influences how children interact with space by assisting them to recognize the importance of landmarks in identifying locations (Shusterman et al., 2011). Studying VLN not only enhances the development of embodied AI that follows human instructions in visual environments, but also deepens our understanding of how cognitive agents develop navigation skills, adapt to different environments, and how language use is connected to visual perceptions and actions.
2.2 Relevant Tasks and Scope of the Survey
Following natural language navigation instructions has traditionally been modeled using symbolic world representations such as maps (Anderson et al., 1991; MacMahon et al., 2006; Paz-Argaman and Tsarfaty, 2019). However, our survey focuses on models that employ visual environments and address the challenges of multimodal understanding and grounding. Likewise, we redirect readers to extensive surveys on visual navigation (Zhu et al., 2021b; Zhang et al., 2022a; Zhu et al., 2022) and mobile robot navigation (Gul et al., 2019; Crespo et al., 2020; Möller et al., 2021), which concentrate on visual perception and physical embodiment. However, these studies provide minimal discussions on the role of language in navigation tasks. While we inevitably extend our discussions of VLN to encompass areas beyond navigation, such as mobile manipulation and dialogue, our primary focus remains on navigational tasks, for which we provide a detailed literature review. Besides, unlike previous VLN surveys (Gu et al., 2022; Park and Kim, 2023; Wu et al., 2024), which offer a bottom-up summary focusing on benchmarks and modeling innovations, our survey adopts a top-down approach, and uses the roles of foundation models to categorize the research efforts into three fundamental challenges from the aspects of the world model, the human model, and the VLN agent. Note that this survey concentrates on frontier methods associated with the rise of foundation models. Thus, we point to less relevant methods (e.g., LSTM-based methods) very briefly at the beginning of each section to motivate our discussions.
Figure 2: VLN challenges and solutions within the framework of world model, human model, and VLN agent. We discuss history and memory in the world model, ambiguous instructions in the human model, generalization ability in them both. For the VLN agent, we discuss methods for grounding and reasoning, planning, and adapting foundation models as agents. Depending on the role served by the foundation models, we categorize these methods into four types. Additionally, we discuss the potential future of the foundation model for the VLN task.
2.3 Task Formulations and Benchmarks
VLN Task Definition.
A typical VLN agent receives a (sequence of) language instruction(s) from human instructors at a designated position. The agent navigates through the environment using an egocentric visual perspective. By following the instructions, its task is to generate a trajectory over a sequence of discrete views or lower level actions and control (e.g., FORWARD $0.25$ meter) to reach the destination, which is considered successful if the agent arrives within a specified distance (e.g., $3$ meters) from the destination. Besides, the agent may exchange information with the instructor during navigation, either by requesting help or engaging in freeform language communication. Additionally, there has been an increasing expectation for VLN agents to integrate additional tasks such as manipulation (Shridhar et al., 2020) and object detection (Qi et al., 2020b), along with navigation.
Benchmarks.
As is summarized in Table 1, existing VLN benchmarks can be taxonomized based on several key aspects: (1) the world where navigation occurs, including the domain (indoors or outdoors) and the specifics of the environment. (2) the type of human interaction involved, including the interaction turns (single or multiple), communication format (freeform dialogue, restricted dialogue, or multiple instructions), and language granularity (action-directed or goal-directed). (3) the VLN agent, including its types (e.g., household robots, autonomous driving vehicles, or autonomous aerial vehicles), action space (graph-based, discrete, or continuous), and additional tasks (manipulation and object detection). (4) the dataset collection, including text collection method (human-generated or templated) and route demonstrations (human-performed or planner-generated). Representatively, Anderson et al. (2018) create the Room-to-Room (R2R) dataset based on the Matterport3D simulator (Chang et al., 2017), where an agent needs to follow fine-grained navigation instructions to reach the goal. Room-across-Room (RxR) (Ku et al., 2020) is a multilingual variation, including English, Hindi, and Telugu instructions. It offers a larger sample size and provides time-aligned instructions for virtual poses, enriching the task’s linguistic and spatial information. Matterport3D allows VLN agents to operate in a discrete environment and rely on pre-defined connectivity graphs for navigation, where agents travel on the graph by teleportation between adjacent nodes, referred to as VLN-DE. To make the simplified setting more realistic, Krantz et al. (2020); Li et al. (2022c); Irshad et al. (2021) propose VLN in continuous environments (VLN-CE) by transferring discrete R2R paths to continuous spaces (Savva et al., 2019). Robo-VLN (Irshad et al., 2021) further narrows the sim-to-real gap by introducing VLN with continuous action spaces that are more realistic in robotics settings. Recent VLN benchmarks have undergone several design changes and expectations, which we discuss in § 6.
Evaluation Metrics.
Three main metrics have been employed to evaluate navigation wayfinding performance (Anderson et al., 2018): (1) Navigation Error (NE), the mean of the shortest path distance between the agent’s final position and the goal destination; (2) Success Rate (SR), the percentage of the final position being close enough to the goal destination; and (3) Success Rate Weighted Path Length (SPL), which normalizes success rate by trajectory length. Some other metrics are used to measure the faithfulness of instruction following and the fidelity between the predicted and the ground-truth trajectory, for example: (4) Coverage Weighted by Length Score (CLS) (Jain et al., 2019); (5) Normalized Dynamic Time Warping (nDTW) (Ilharco et al., 2019), which penalizes deviations from the ground-truth trajectories; and (6) Normalized Dynamic Time Warping Weighted by Success Rate (sDTW) (Ilharco et al., 2019), which penalizes deviations from the ground-truth trajectories and also considers the success rate.
3 World Model: Learning and Representing the Visual Environments
A world model helps the VLN agent to understand their surrounding environments, predict how their actions would change the world state, and align their perception and actions with language instructions. Two challenges have been highlighted in existing work about learning a world model: encoding the visual observation history in the current episode in the memory, and generalization to unseen environments in new tasks.
3.1 History and Memory
Different from other vision-language tasks like Visual Question Answering (VQA) (Antol et al., 2015), Visual Entailment (Xie et al., 2019), etc, the VLN agent needs to incorporate the history information of past actions and observations into its current step’s input to determine the action rather than solely consider image and text in a single step. Prior to employing the foundation models in VLN, LSTM hidden states serve as an implicit memory supporting agents’ decision-making during navigation, and researchers further design different attention mechanisms (Tan et al., 2019; Wang et al., 2019) or auxiliary tasks (Ma et al., 2019; Zhu et al., 2020) to improve the alignment between the encoded history and instructions.
History Encoding.
With the advance in foundation models, many works build the VLN agent with Transformer architecture and propose different ways to explicitly encode the navigation history. Hong et al. (2021) first proposes to utilize the [CLS] token from the transformer recurrently for encoding the history information. Pashevich et al. (2021) explicitly encodes the history observations with a multi-modal Transformer. Chen et al. (2021b) further improves the history encoding with a panorama encoder to encode the visual observation at each time step, and a history encoder to encode all the past observations. This design first eliminates the need of recurrent state for history encoding, and enables efficient and large-scale pre-training on instruction-path pairs. Follow-up research replaces the panorama encoder with mean pooling of images (Kamath et al., 2023) or front-view image encoding (Qiao et al., 2022). Lin et al. (2022a) further proposes a variable-length memory framework that stores the previous activations in a memory bank and introduces the memory-aware consistency loss to learn the relations between instruction and temporal context in the memory bank.
Graph-based History.
Another line of research enhances the navigation history modeling with graph information. For example, some works utilize a structured Transformer encoder to capture the geometric cues in the environment (Chen et al., 2022d; Deng et al., 2020; Wang et al., 2023b; Zhou and Mu, 2023; Su et al., 2023; Zheng et al., 2023; Wang et al., 2021; Chen et al., 2021a; Zhu et al., 2021a). In addition to the topological graph used in encoding, many works propose to include the top-down view information (e.g., grid map (Wang et al., 2023e; Liu et al., 2023a), semantic map (Hong et al., 2023a; Huang et al., 2023; Georgakis et al., 2022; Anderson et al., 2019; Chen et al., 2022b; Irshad et al., 2022), local metrics map (An et al., 2023)), and local neighborhood map (Gopinathan et al., 2023) in modeling the observation history during navigation.
3.2 Generalization across Environments
One main challenge in the VLN is learning from limited available environments and generalizing to new and unseen environments. Many works demonstrate that learning from semantic segmentation features (Zhang et al., 2020), dropout information in the environment during training (Tan et al., 2019), and maximizing the similarity between semantically-aligned image pairs from different environments (Li et al., 2022a) improve agents’ generalization performance to unseen environments. These observations suggest the need to learn from large-scale environment data to avoid overfitting to training environments. Next, we discuss how existing works collect new environment data, and utilize it in training.
Pre-trained Visual Representations.
Most works obtain vision representations from ResNet pre-trained on ImageNet (Anderson et al., 2018; Tan et al., 2019). Shen et al. (2021) replace it with the CLIP visual encoder (Radford et al., 2021), which is pre-trained with contrastive loss between image-text pairs and naturally aligns better with the instructions, boosting the VLN performance. Wang et al. (2022b) further explores transferring vision representation learned from video data for VLN task, suggesting that temporal information learned from video is crucial for navigation.
Environment Augmentation.
One main line of research focuses on augmenting the navigation environment with auto-generated synthetic data. EnvEdit (Li et al., 2022b), EnvMix (Liu et al., 2021), KED (Zhu et al., 2023), and FDA (He et al., 2024a) generate synthetic data by changing the existing environments from Matterport3D. Specifically, they mix up rooms from different environments, change appearance and style of the environments, and interpolate high-frequency features with the environments. Pathdreamer (Koh et al., 2021) and SE3DS (Koh et al., 2023) further synthesize the environments in future steps given current observations and explore utilizing the synthesis view as augmented data for VLN training.
Different from all these approaches which generate new environments mostly conditioning on the existing environments, another line of research introduces new environments from different domains for VLN learning. AirBert (Guhur et al., 2021) introduces a pipeline to transferring single images from Airbnb to panorama observations in the navigation trajectory. Similarly, Lily (Lin et al., 2023b; Li et al., 2024a) introduces videos and sample trajectories from frames in the videos as augmentation data. While there still exists some domain gap between Airbnb images, Youtube videos and actual navigation environments, ScaleVLN (Wang et al., 2023f), HM3DAutoVLN (Chen et al., 2022c), and MARVAL (Kamath et al., 2023) collect trajectory-instruction pairs from un-annotated HM3D (Ramakrishnan et al., 2021) and Gibson (Xia et al., 2018) environments. The agent fine-tuned on data collected on these environments shows strong generalization performance when given under-specified instructions (Qi et al., 2020b), long instructions (Ku et al., 2020), and step-by-step instruction (Anderson et al., 2018). PanoGen (Li and Bansal, 2024) further explores synthesizing new panorama environments conditioned on text, which shows the potential of collecting large-scale panorama environments without any existing 3D simulation environments.
The learning paradigm from the collected environments has changed with the advances in foundation models. Prior to the prevalence of pre-training in foundation models, most works directly augment the training environment with the auto-collected new environments during VLN fine-tuning (Li et al., 2022b; Liu et al., 2021; Koh et al., 2021, 2023; Zhu et al., 2023). As pre-training has been demonstrated to be crucial for foundation models, it has also become a standard practice in VLN to learn from collected environments during the pre-training stage (Li and Bansal, 2024; Kamath et al., 2023; Chen et al., 2022c; Wang et al., 2023f; Lin et al., 2023b; Guhur et al., 2021; He et al., 2024a). The large-scale pre-training with the augmented in-domain data has become crucial in bridging the gap between the performance of agent and human.
4 Human Model: Interpreting and Communicating with Humans
Besides learning and modeling the world, VLN agents also need a human model that comprehends human-provided natural language instructions per situation to complete navigation tasks. There are two main challenges: resolving ambiguity and generalization of grounded instructions in different visual environments.
4.1 Ambiguous Instructions
Ambiguous instructions mainly arise in single-turn navigation scenarios, where the agent follows an initial instruction without further human interaction for clarification. These instructions lack the flexibility to train the agent to adapt its language understanding and visual perception to the dynamic environment. For instance, instructions may contain landmarks invisible at the current view or indistinguishable landmarks visible from multiple views Zhang and Kordjamshidi (2023).. The issue of ambiguous instructions is barely addressed before the application of foundational models to VLN. Although LEO Xia et al. (2020) attempts to aggregate multiple instructions to describe the same trajectory from different perspectives, it still relies on human-annotated instructions. However, comprehensive perceptual context and commonsense knowledge from foundational models enable the agent to interpret ambiguous instructions using external knowledge, as well as seek assistance from other human models.
Perceptual Context and Commonsense Knowledge.
Large-scale cross-modal pre-trained models like CLIP are capable of matching visual semantics with text. This enables the VLN agent to utilize information from the visual objects and their states in the current perception to resolve ambiguity, especially in single-turn navigation scenarios. For example, VLN-Trans (Zhang and Kordjamshidi, 2023) constructs easy-to-follow sub-instructions with visible and distinctive objects obtained from CLIP to pre-train a Translator that converts original ambiguous instructions into easily understandable sub-instruction representations. LANA+ (Wang et al., 2023d) leverages CLIP to query a text list of landmark semantic tags with the visual panoramic observations, and selects the top-ranked retrieved textual cues as representations of the salient landmarks to follow. NavHint (Zhang et al., 2024b) constructs a hint dataset, providing detailed visual descriptions to help the VLN agent build a comprehensive understanding of the visual environment rather than focusing solely on the objects mentioned in the instructions. On the other hand, the commonsense reasoning ability of LLMs can be used to clarify or correct ambiguous landmarks in the instructions, and break instructions into actionable items. For example, Lin et al. (2024b) use LLMs to provide commonsense about open-world landmark co-occurrences and conduct CLIP-driven landmark discovery accordingly. SayCan (Ahn et al., 2022) breaks an instruction into a ranked list of pre-defined admissible actions and combines them with an affordance function that assigns higher weights to the objects appearing in the current scene.
Information Seeking.
While ambiguous instructions can be resolved based on visual perception and situational context, another more direct approach is to seek help from the communication partner, i.e., the human speakers who generate the instructions (Nguyen and Daumé III, 2019; Paul et al., 2022). There are three key challenges in this line of work: (1) deciding when to ask for help (Chi et al., 2020); (2) generating information-seeking questions, e.g., next action, objects, and directions (Roman et al., 2020; Singh et al., 2022); (3) developing an oracle that provides the queried information, which could be either real humans (Singh et al., 2022), rules and templates (Gao et al., 2022), or neural models (Nguyen and Daumé III, 2019). LLMs and VLMs could potentially fit two roles in this framework, either as an information-seeking model or information-providing model. Fan et al. (2023b) use GPT-3 to preprocess the ground-truth responses in the training data in a step-by-step manner and train an oracle using a pre-trained SwinBert (Lin et al., 2022b) video-language model. They also demonstrate off-the-shelf large vision-language models like mPLUG-Owl (Ye et al., 2023) can serve as strong zero-shot oracles. Besides, self-motivated communication agents have been proposed (Zhu et al., 2021c) by learning the confidence of the oracle to produce a positive answer, which enables self-Q&A manner where the oracle can be removed at inference time.
4.2 Generalization of Grounded Instructions
The limited scale and diversity of navigation data is another significant issue affecting the VLN agent’s ability to comprehend various linguistic expressions and follow instructions effectively, particularly in unseen navigation environments. Although the language style itself has good generalization capability across seen and unseen environments Zhang et al. (2020), how to ground the instructions with the unseen environments is potentially a hard task given the limited scale of training instructions. Foundation models help address these issues through both pre-trained representations and instruction generation for data augmentation.
Pre-trained Text Representations.
Before the foundation models, many works rely on text encoders, such as LSTM, to represent text instructions Anderson et al. (2018); Tan et al. (2019). The foundation models significantly enhance the VLN agent’s language generalization ability through pre-trained representations. For example, PRESS Li et al. (2019) fine-tunes the pre-trained language model BERT Kenton and Toutanova (2019) to obtain text representations that generalize better to previously unseen instructions. The multi-modal Transformers Tan and Bansal (2019); Lu et al. (2019) boost methods, such as VLN-BERT Majumdar et al. (2020) and PREVALENT Hao et al. (2020), to obtain more generic vision-linguistic representations by pre-training on large-scale text-image pairs collected from the web.
Instruction Synthesis.
Another method to improve the agent’s generalization ability is to synthesize more instructions. Early works employ the Speaker-Follower framework Fried et al. (2018); Tan et al. (2019); Kurita and Cho (2020); Guhur et al. (2021) to train an offline speaker (instruction generator) using human-annotated instruction-trajectory pairs. It then generates new instructions based on sequences of panoramas along a given trajectory. However, Zhao et al. (2021) observe that these generated instructions are low-quality and show a poor performance in human wayfinding evaluation. Marky Wang et al. (2022a); Kamath et al. (2023) addresses this limitation using a multi-modal extension of the multilingual T5 model Xue et al. (2020) with text-aligned visual landmark correspondences, achieving near-human quality on R2R-style paths in unseen environments. Additionally, instead of training an instruction generator, some recent research automatically generates instructions using cross-modal Transformer Liang et al. (2022); Lin et al. (2023b); Zhang and Kordjamshidi (2023). For instance, ProbES (Liang et al., 2022) self-explores environments by sampling trajectories and automatically constructs the corresponding instruction by filling the instruction templates with movements and object phrases detected by CLIP.
5 VLN Agent: Learning an Embodied Agent for Reasoning and Planning
While the world and human models empower visual and language understanding abilities, VLN agents need to develop embodied reasoning and planning capabilities to support their decision-making. From this perspective, we discuss two challenges: grounding and reasoning, and planning. We also explore the method of directly applying foundation models as the VLN agent backbone.
5.1 Grounding and Reasoning
Different from other VL tasks, such as VQA and Image Captioning, which primarily focus on static alignment between images and corresponding textual descriptions, the VLN agent needs to reason about spatial and temporal dynamics in the instructions and the environment based on its actions. Specifically, the agent should consider previous actions, identify the part of the sub-instruction to execute, and ground the text to the visual environment to execute the action accordingly. Previous methods primarily rely on explicit semantic modeling or auxiliary task design to obtain such abilities. However, pre-training with specially designed tasks has become the dominant approach with the advent of foundation models.
Explicit Semantic Grounding.
The previous efforts enhance the agent’s grounding ability through explicit semantic modeling in both vision and language modalities, including modeling motions and landmarks (Hong et al., 2020b, a; Zhang et al., 2021; Qi et al., 2020a), utilizing syntactic information in the instruction (Li et al., 2021), as well as spatial relations (Zhang and Kordjamshidi, 2022a). Very few works (Lin et al., 2023a; Zhan et al., 2024a; Wang et al., 2023b) explore explicit grounding in the VLN agent with the foundation models. Lin et al. (2023a) proposes actional atomic-concept learning and map visual observations to actional atomic concepts to facilitate multi-modal alignments.
Pre-training VLN Foundation Models.
Except for explicit semantic modeling, the previous research also enhances the agent’s grounding ability through auxiliary reasoning tasks (Ma et al., 2019; Wu et al., 2021; Zhu et al., 2020; Raychaudhuri et al., 2021; Dou and Peng, 2022; Kim et al., 2021). Such methods are less explored in VLN agents with foundation models, as their pre-training already provides a general understanding of spatial and temporal semantics prior to navigation. Various pre-training methods with specially designed tasks have been proposed to improve the agent’s grounding ability. Lin et al. (2021) introduce pre-training tasks specifically designed for scene and object grounding. LOViS (Zhang and Kordjamshidi, 2022b) formulates two specialized pre-training tasks to enhance orientation and visual information separately. HOP (Qiao et al., 2022, 2023a) introduces a history-and-order aware pre-training paradigm that emphasizes historical information and trajectory orders. Li and Bansal (2023) suggests that enhancing the agent with the ability to predict future view semantics helps the agent in longer path navigation performance. Dou et al. (2023) design a masked path modeling objective to reconstruct the original path given a randomly masked sub-path. Cui et al. (2023) propose entity-aware pre-training by predicting grounded entities and aligning them to text.
5.2 Planning
Dynamic planning enables VLN agents to adapt to environmental changes and improve navigation strategies on the fly. Alongside the graph-based planners that utilize global graph information to enhance local action spaces, the rise of foundational models, particularly LLMs, has brought LLM-based planners into the VLN field. These planners use LLMs’ vast commonsense knowledge and advanced reasoning to create dynamic plans that improve decision-making.
Graph-based Planner.
Recent advancements in VLN emphasize enhancing navigational agents’ planning capabilities through global graph information. Among them, Wang et al. (2021); Chen et al. (2022d); Deng et al. (2020); Zheng et al. (2023) enhance the local navigation action spaces with global action steps from graph frontiers of visited nodes for better global planning. Gao et al. (2023) further enhances navigation decision-making with high-level planning for zone selection and low-level planning for node selection. Moreover, Liu et al. (2023a) enriches the graph-frontier-based global and local action spaces with grid-level actions for more accurate action prediction. In continuous environments, Krantz et al. (2021); Hong et al. (2022); Anderson et al. (2021) adopt a hierarchical planning approach utilizing high-level action spaces instead of low-level ones by selecting a local waypoint from a predicted local navigability graph. CM2 (Georgakis et al., 2022) facilitates trajectory planning by grounding instructions within a local map. Expanding on this strategy, An et al. (2024, 2023); Wang et al. (2023e); Chang et al. (2023b); Wang et al. (2022c) construct a global topological graph or grid maps to facilitate map-based global planning. Additionally, Wang et al. (2023a, 2024a) predict multiple future waypoints using either a video prediction model or a neural radiance representation model to plan the best action based on the long-term effects of predicted candidate waypoints.
LLM-based Planner.
In parallel, some studies leverage common-sense knowledge from LLMs to generate text-based plans (Huang et al., 2022a, b). LLM-Planner (Song et al., 2022) creates detailed plans composed of sub-goals, dynamically adjusting these plans in real-time by integrating detected objects according to predefined program patterns. Similarly, Mic (Qiao et al., 2023b) and $A^{2}$av (Chen et al., 2023b) specialize in breaking down navigation tasks into detailed textual instructions, with Mic generating step-by-step plans from both static and dynamic perspectives, while $A^{2}$av uses GPT-3 to parse instructions into actionable sub-tasks. ThinkBot (Lu et al., 2023) employs thought chain reasoning to generate missing actions with interactive objects. Additionally, SayNav (Rajvanshi et al., 2023) builds a 3D scene graph of the explored environment as input to LLMs for generating feasible and contextually appropriate high-level plans for the navigator.
5.3 Foundation Models as VLN Agents
The architecture of VLN agents has undergone significant transformations with the advent of foundation models. Initially conceptualized by Anderson et al. (2018), VLN agents were formulated within a Seq2Seq framework, employing an LSTM and an attention mechanism to model the interaction between vision and language modalities. With the advent of foundation models, the agent backend has transitioned from LSTM to Transformer and, more recently, to these large-scale pre-trained systems.
VLMs as Agents.
The mainstream methodology leverages single-stream VLMs as the core structure of VLN agents (Hong et al., 2021; Qi et al., 2021; Moudgil et al., 2021; Zhao et al., 2022). These models process inputs from language, vision, and historical tokens simultaneously at each time step. It performs self-attention over these cross-modal tokens to capture the textual-visual correspondence, which is then used to infer the action probability. VLN-CE agents (Krantz et al., 2020) differentiate themselves from the VLN-DE (Anderson et al., 2018) agents by their action space, executing low-level controls in the continuous environment instead of graph-based high-level actions of view selection. Despite early works (Krantz et al., 2020; Raychaudhuri et al., 2021) utilizing LSTM to infer low-level actions, the introduction of waypoint predictors has allowed to transfer methods from DE to CE (Krantz et al., 2021; Krantz and Lee, 2022; Hong et al., 2022; Anderson et al., 2021) All these methods use a waypoint predictor to get a local navigability graph, allowing foundation models in DE to adapt to the continuous environment.
LLMs as Agents.
Since LLMs have powerful reasoning ability and semantic abstraction of the world, and also show strong generalization ability in unknown large-scale environments, recent research in VLN has started to directly employ LLMs as agents to complete navigation. Typically, visual observations are converted into textual descriptions and fed into the LLM along with instructions, which then perform action predictions. Innovations such as NavGPT (Zhou et al., 2024a) and MapGPT (Chen et al., 2024a) demonstrate the feasibility of zero-shot navigation, with NavGPT autonomously generating actions using GPT-4 and MapGPT converting topological maps into global exploration hints. DiscussNav (Long et al., 2023) extends this approach by deploying multiple domain-specific LLM experts to automate and reduce human involvement in navigation tasks. Some studies have incorporated the Chain-of-Thought (CoT) (Wei et al., 2022) reasoning mechanism to improve the reasoning process. NavCoT (Lin et al., 2024a) transforms LLMs into a world model and navigational reasoning agent, streamlining decisions by simulating future environments. MC-GPT (Zhan et al., 2024b) employs memory topology maps and human navigation examples to diversify strategies, while InstructNav (Long et al., 2024) breaks navigation into sub-tasks with multi-sourced value maps for effective execution. In contrast to zero-shot usage, some works (Zheng et al., 2024; Zhang et al., 2024a; Pan et al., 2023) fine-tune LLMs to address the embodied navigation tasks effectively. This demonstrates the flexibility and practical potential of fine-tuned language models in both simulation and real-world scenarios, marking a significant advancement over traditional applications.
6 Challenges and Future Directions
While foundation models have enabled novel solutions to VLN, several limitations remain underexplored and new challenges arise. In this section, we delve into the challenges and future direction of the VLN field from the perspectives of benchmarks, the world model, the human model, the agent model, and real robot deployment.
Benchmarks: Limitations of Data and Task.
The current VLN datasets have limitations regarding quality, diversity, bias, and scalability. For example, in the R2R dataset, the instruction-trajectory pairs are biased to the shortest path, which may not accurately represent real-world navigation scenarios. We discuss the trends and recommendations how VLN benchmarks can be improved.
  • Unified Tasks and Platforms. Establishing robust benchmarks and ensuring reproducibility are crucial for evaluating VLN in real-world settings. Real-world variability necessitates comprehensive benchmarks reflecting navigation challenges. A universal sim-to-real evaluation platform, like OVMM (Yenamandra et al., 2024), is needed for standardized testing across simulated and real-world settings.
  • Dynamic Environment. Real-world environments are inherently complex and dynamic, with moving objects, people, and variations like lighting and weather presenting unexpected situations (Ma et al., 2022). These factors disrupt the visual perception of navigation systems and make maintaining reliable performance difficult. Recent efforts like HAZARD (Zhou et al., 2024b), Habitat 3.0 (Puig et al., 2024), and HA-VLN (Li et al., 2024b) consider dynamic environments and provide a good starting point.
  • Indoors to Outdoors. VLN agents navigating in outdoor environments, e.g., autonomous driving and aerial vehicles, also start to get more attention (Vasudevan et al., 2021; Li et al., 2024a), with various language-guided datasets (Sriram et al., 2019; Ma et al., 2022) developed. Early studies have attempted to involve LLMs in these tasks, either with prompt engineering (Shah et al., 2023; Sha et al., 2023; Wen et al., 2023), or by fine-tuning LLMs to predict the next action or plan future trajectories (Chen et al., 2023a; Mao et al., 2023). To adapt off-the-shelf VLMs to these outdoor navigation domains, real-world driving videos (Xu et al., 2023; Yuan et al., 2024), simulated driving data (Wang et al., 2023c; Shao et al., 2023) and them both (Sima et al., 2023; Huang et al., 2024b) have been utilized for instruction tuning so that these foundation models learn to predict future throttle and steering angles. Additional reasoning and planning modules have also been integrated into foundation model driving agents (Huang et al., 2024b; Tian et al., 2024). We refer the readers to surveys and position papers for a detailed review (Li et al., 2023a; Cui et al., 2024; Gao et al., 2024; Yan et al., 2024).
World Model: From 2D to 3D.
Building effective world representations is a central research theme in embodied perception, reasoning, and planning. Although the current research represents the world with strong and generic 2D representations, VLN is fundamentally a 3D task, where the agent perceives the real-world environment in 3D. Many explicit 3D representations are developed in prior work, including various semantic SLAMs and volumetric representation (Chaplot et al., 2020; Min et al., 2021; Saha et al., 2022; Blukis et al., 2022; Zhang et al., 2022b; Liu et al., 2024), depth information (An et al., 2023), Bird’s-Eye-View representations like grid map (Wang et al., 2023e; Liu et al., 2023a), and local metrics map (An et al., 2023). These representations are limited because they reduce the object set to a closed set, making them inadequate for open-vocabulary settings with natural language. Several studies develop queryable map/scene representations by integrating multi-view image features captured from CLIP into 3D voxel grids (Jatavallabhula et al., 2023; Chang et al., 2023a) or top-down feature maps (Huang et al., 2023; Chen et al., 2022a), as well as utilizing scene graphs (Rana et al., 2023; Gu et al., 2023b) to represent spatial relationships. However, adapting 3D representations learned from large-scale data for VLN agents to better perceive the 3D environment is still under exploration. The recent rise of 3D foundation models, including 3D reconstruction models (Hong et al., 2024) and 3D multimodal language models (Hong et al., 2023b; Yang et al., 2024; Huang et al., 2024a), can be crucial for VLN.
Human Model: From Instruction to Dialogue.
Previous efforts predominantly adopt either a speaker-listener paradigm or restricted QA dialogue (Thomason et al., 2020; Gao et al., 2022) that only allows the agent to ask for help. Recently, there has been a surge in new benchmarks featuring open-ended dialogue instructions (De Vries et al., 2018; Banerjee et al., 2021; Padmakumar et al., 2022; Ma et al., 2022; Fan et al., 2023a), supporting fully free-form communication where agents can ask, propose, explain, suggest, clarify, and negotiate even in ambiguous or confusing scenarios. Still, current approaches rely on rule-based dialogue templates to tackle these complexities (Zhang et al., 2023; Parekh et al., 2023; Gu et al., 2023a), though they might feature a foundation model component. Huang et al. (2024b) perform conversational tuning on a video-language model using human-human dialogue data paired with simulated navigation videos, showcasing enhanced dialogue generation capabilities while navigation. Moving forward, it is imperative for future research to integrate foundation models for situated task-oriented dialogue management (Ulmer et al., 2024), or explore existing foundation models for task-oriented dialogue (He et al., 2022).
Agent Model: Adapting Foundation Models for VLN.
While foundation models show strong generalizability, incorporating them into navigation tasks remains challenging. LLMs fundamentally lack the capability to visually perceive the actual environment and are prone to hallucinations.
  • Lack of Embodied Experience. This limitation can lead to scenarios where LLMs rely solely on pre-established commonsense for task planning and reasoning, which might not meet specific real-world needs (Xiang et al., 2024). Some pipelines tackle this issue by captioning the visual observations to textual descriptions as prompts for LLMs (Zheng et al., 2022), with a potential loss of essential visual semantics. Compared with LLMs, VLM agents demonstrate the potential to perceive the visual world and plan (Zhang et al., 2024a). Still, these models are primarily developed from internet data, which lack embodied experiences (Mu et al., 2024) and need finetuning for robust agentic decision-making (Zhai et al., 2024). Further research is needed to transfer the commonsense knowledge in foundation model agents to generalize to embodied situations.
  • Hallucination Issue. LLMs and VLMs might generate non-existent objects, leading to misinformation (Li et al., 2023b; Chen et al., 2024b). For example, when LLM performs task planning, it may generate instructions such as “go forward and turn left at the sofa” even if there is no sofa in the room. This inaccuracy may cause them to execute incorrect or impossible actions.
Deployment: From Simulation to Real Robots.
Simulated settings often lack the complexity and variability of real-world environments, and lower-quality rendered images exacerbate this issue. First, the perception gap results in decreased performance and accuracy, highlighting the need for more robust perception systems. Wang et al. (2024b) have started to explore the use of semantic maps and 3D feature fields to provide monocular robots with panoramic perception shows improved performance. The embodiment gap and the data scarcity are also bottlenecks. The rise of robot teleportation (He et al., 2024b) also provides an alternative to scale up VLN data for foundation models in real human-robot communications.
Acknowledgement
This work is supported in part by the ARO Award W911NF2110220, NSF grant IIS-1949634, and ONR grant N00014-23-1-2417 & N00014-23-1-2356. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funding agencies.